# Distributed Motion Control of Multiple Mobile Manipulators for Reducing Interaction Wrench in Object Manipulation

### v2

@@ III-C Simulations
TABLE I: The actual end velocity (m/s) of each robot
@@
Due to some uncertainties and the deviation between the planned robot model and the actual robot model, it is impossible to achieve the planned velocity. The actual end velocities of robots are shown in Table I. If the deviation is not corrected by employing the control law, it will result in increased interaction wrenches, as illustrated in Fig. 2(a). In the simulation, two types of errors are introduced. One type is random velocity error, and the other is velocity bias. The bias can be caused by the relative pose errors of the robots. Both types of errors are very common in practice, and may cause increased interaction wrenches.
@@
The proposed control law Eq.(9) is verified by enrolling it after 20s in the process described above. The communication topology is shown in Fig. 3. As described before, time delays are non-uniform, asymmetric, and time-varying, and they are assumed to be $\tau_{12}=0.01||sin(t)||$, $\tau_{14}=0.01||cos(t)||$, $\tau_{21}=0.02/t$, $\tau_{23}=0.02/t^{2}$, $\tau_{24}=e^{-t}$, $\tau_{25}=t/e^{t}$, $\tau_{34}=0.004In(t)$, $\tau_{35}=0.02/In(t)$, $\tau_{45}=0.01||\Omega(t)||$, $\tau_{51}=0.01||\Omega(t)||$. The control coefficients are set as $K_{0}=\text{diag}\left\{10,10\right\}$ and $k=0.5,\beta=0.1$. According to these time delays and $t\in[20,60]$, the bound $\tau$ is 0.01s. These parameters substituted into Eq.(28) satisfy $\alpha_{1},\alpha_{2},\alpha_{3}>0$. Therefore, the control law Eq.(9) is stable. The results are illustrated in Fig. 2(b). After 20 seconds, the control law is enrolled, and the force data is quickly adjusted to the desired value. Although we use five robots as an example, the control law can easily be extended to more robots. The computation of the control law Eq.(9) itself is not complex.
Figure 4: The physical model of a mobile manipulator robot used in experiments. The robot consists of an omnidirectional wheeled platform and a 6-DOF manipulator. A force sensor and a gripper are equipped at the end of the robot for manipulating objects. A router is attached for communication between different robots.
@@
Figure 6: Comparison between applying and not applying the control law. The interaction wrenches in the world frame without (with) applying the control law is represented by the red (blue) curve. The desired interaction wrenches are $[10,0,-7,0,0,0]^{\text{T}}$ for robot1 and $[-10,0,-7,0,0,0]^{\text{T}}$ for robot2.
Figure 7: The total wrenches applied to the object during cooperation manipulation by multiple mobile manipulator robots.
IV Experiments and Discussions
IV-A Experimental Setup
The proposed control law is verified in physical experiments. Each robot is composed of a HANGFA omnidirectional wheeled platform and a JAKA Zu7 manipulator. The end of each manipulator is equipped with a Robotiq FT 300-S Force Torque Sensor and a Robotiq 2F-85 Gripper. These devices and the robot are controlled by its onboard Next Unit of Computing (NUC) computer with Intel Core i7-1165G7 CPU at 2.80 GHz (4 cores) and 16GB RAM. The control frequency of the robots is 25 Hz. Besides, each robot is equipped with wireless transmission modules in the same local network for communication. The communication framework is implemented and achieved through ROS-melodic. Generally, the communication delays are less than 0.01s. The physical robot is shown in Fig. 4. The forward and inverse kinematics of the robot are based on our previous work [36].
Figure 8: Two mobile manipulator robots cooperatively transport an object. They first lift the object, then move it to the target position, and finally put it down.
Figure 9: Comparison between applying and not applying the control law. The force sensor data without (with) applying the control law is represented by the red (blue) curve. The desired interaction wrenches are $[10,0,-5,0,0,0]^{\text{T}}$ for robot1 and $[-10,0,-5,0,0,0]^{\text{T}}$ for robot2.
Figure 10: The total force applied to the object during cooperation manipulation by multiple mobile manipulator robots.
IV-B Experimental Results
@@
The process of the experiment is illustrated in Fig. 5, and the motion duration is 83.2 seconds. The experiment is repeated twice, once without the control law and once with. The results of interaction wrenches are shown in Fig. 6. Without the control law, the interaction wrenches could not be maintained at the desired values. Especially in the $x$-direction, the forces are very large due to the pushing and pulling between the robots when cooperatively manipulating the object. Additionally, in the $z$-direction, the robots could not evenly share the load. These increased interaction wrenches do not contribute to supporting or manipulating the object, instead, they may cause damage to both the object and the robots. In contrast, with the control law applied, the interaction wrenches are maintained at the desired values in all six directions. Compared to the scenario without the control law, implementing the control law reduces the maximum error by [93.37, 66.62, 92.38, 68.75, 58.82, 64.24]and the average error by [98.68, 93.49, 96.58, 55.36, 61.96, 84.89]for robot1 in the corresponding directions, and the maximum error by [94.38, 72.55, 93.30, 75.74, 70.18, 26.05]and the average error by [98.56, 90.15, 97.30, 74.94, 73.36, 37.10]for robot2 in the corresponding directions. The total wrenches on the object are shown in Fig. 7. It can be seen that after applying the control law, the fluctuation range of the total wrenches on the object is also relatively smaller.
@@
The process of the experiment is illustrated in Fig. 8, and the motion duration is 130.4 seconds. From 17s to 57s, the platforms move forward simultaneously to avoid exceeding the reachable space of the manipulators. From 70s to 100s, the platforms move laterally simultaneously to transport the object. As shown in Fig. 9, it can be seen that without the control law, the interaction wrenches vary greatly, especially in the $f_{x}$ directions. Compared to the scenario without the control law, implementing the control law reduces the maximum error by [45.97, 54.18, 53.56, 81.85, 77.39, 21.52]and the average error by [85.92, 62.72, 60.49, 93.13, 88.80, 67.99]for robot1 in the corresponding directions, and the maximum error by [45.47, 34.97, 52.53, 73.38, 74.41, 46.05]and the average error by [86.38, 58.44, 69.38, 85.51, 83.32, 72.44]for robot2 in the corresponding directions.
@@
IV-C Discussion and Limitation
The stability of the proposed control law has been proven, and subsequent simulations and experiments have demonstrated that the proposed control law effectively maintains the desired interaction wrenches. Due to positioning errors or control errors, the theoretical planned motion cannot achieve desired cooperative performance, and the errors will result in increased internal wrenches. Reducing unnecessary interaction wrenches is essential for ensuring the safety of both the object and the robots, and reducing emergency stops during collaborative manipulation. Compared to other control methods, the proposed control law only utilizes local force sensor information, eliminating the need for a large range of accurate positioning devices. Therefore, it is more suitable for real-world applications. Additionally, the communication delay between robots are considered, and it has been proven that the control law is robust against delays.
However, the proposed control strategy also has some limitations. First, the stiffness of the object is required. This parameter often requires empirical values or preliminary experimental determination. Improper setting may lead to instability or divergence of the control law. Second, the effect of acceleration is not considered in Eq.(8) The control law is only suitable for quasi-static processes, so the impact in the experiment cannot be corrected, as shown in Fig. 9. Finally, the motion of the mobile platform has not been taken into account to correct errors, which could lead to failure if the error becomes too large and exceeds the reachable space of the manipulators.
V Conclusion and Future Work
This paper aims to reduce the interaction wrenches during the cooperative object manipulation of multiple mobile manipulator robots. A distributed motion control law is proposed. Since the manipulator is more accurate than the mobile platform, the control law only operates on the manipulators. The control law is proven to be stable under certain conditions. Then, the control law is verified in real-world experiments. Experimental results, involving the cooperative manipulation of two robots transporting an object, effectively validate the effectiveness of the proposed control law. It is shown that the proposed control law is convenient to be applied in real-world applications as it does not rely on global position and velocity information.
Regarding future work, on the one hand, we will further refine the control strategy. The dynamics model of the robot will be considered to reduce impact forces when the robot’s velocity changes. On the other hand, more complex manipulation tasks and manipulating heavier objects will be considered. We will strive to further enhance the control method through real-world manipulation tasks.

### v3

@@ 3.3 Simulations
Due to some uncertainties and the deviation between the planned robot model and the actual robot model, it is impossible to achieve the planned velocity. The actual end velocities of robots are shown in Table 2. If the deviation is not corrected by employing the control law, it will result in increased interaction wrenches, as illustrated in Fig. 4(a). In the simulation, two types of errors are introduced. One type is random velocity error, and the other is velocity bias. The bias can be caused by the relative pose errors of the robots. Both types of errors are very common in practice, and may cause increased interaction wrenches.
Table 2: The actual end velocity (m/s) of each robot
@@
Figure 4: The interaction wrenches between robots and the object. (a) Without the proposed control law. (b)-(d) With the proposed control law.
@@
The proposed control law Eq.(9) is verified by enrolling it after 40s in the process described above. The communication topology is shown in Fig. 3. As described before, time delays are non-uniform, asymmetric, and time-varying, and they are assumed to be $\tau_{12}=||0.01\text{sin}(t)||,\tau_{13}=||0.01\Omega(t)||,\tau_{14}=||0.01\text{cos}(t)||,\tau_{15}=||0.01\Omega(t)||,\tau_{21}=\frac{0.02}{t},\tau_{23}=\frac{0.02}{t^{2}},\tau_{24}=\text{e}^{-t},\tau_{25}=t\text{e}^{-t},\tau_{31}=||0.01\Omega(t)||,\tau_{32}=||0.01\Omega(t)||,\tau_{34}=0.002\text{ln}t,\tau_{35}=\frac{0.02}{\text{ln}t},\tau_{41}=\frac{0.02}{\text{ln}t},\tau_{42}=||0.01\text{sin}(t)||,\tau_{43}=||0.01\text{cos}(t)||,\tau_{45}=||0.01\Omega(t)||,\tau_{51}=||0.01\Omega(t)||,\tau_{52}=||0.01\Omega(t)||,\tau_{53}=||0.01\Omega(t)||,\tau_{54}=||0.01\Omega(t)||$. The control coefficients are set as $k=0.5,\beta=0.1$. The control frequency of the robots is 25 Hz. According to these time delays and $t\in[40,60]$, the bound $\tau$ is 0.01s. These parameters substituted into Eq.(28) satisfy $\alpha_{1},\alpha_{2},\alpha_{3}>0$. Therefore, the control law Eq.(9) is stable. The result for topology Fig. 3(a) is illustrated in Fig. 4(b). After 40s, the control law is enrolled, and the wrench is quickly adjusted to the desired value. Although we use five robots as an example, the control law can easily be extended to more robots. The computation of the control law Eq.(9) itself is not complex.
To test the effect of communication graphs with different connectivity on the control law, the graph in Fig. 3(b) is used. The parameters and other conditions are identical, and the result is shown in Fig. 4(c). If stability is defined as the errors being less than 0.5N, Fig. 4(b) requires 1.12s after applying the control law, while Fig. 4(c) requires 0.32s. Since Fig. 3(b) is a complete graph with stronger connectivity than Fig. 3(a), the robots can communicate and share information more directly and efficiently. Therefore, the convergence speed is much faster. For the proposed distributed control law, when the parameters are the same, stronger graph connectivity leads to faster convergence.
Figure 5: The effect of the delay upper bound on the control law.
To test the effect of communication delay on the control law, the upper bound of the delay is adjusted to 5s. Other parameters and conditions are the same with those in Fig. 4(b). The result is shown in Fig. 4(d). After the control law is applied, it takes 7.84s to reach stability. To further investigate the impact of time delays, tests are conducted with various delay upper bounds. Since both delays and disturbances are random errors, each test is repeated 100 times, and the average values for convergence speed and steady-state errors are recorded. The results are shown in Fig. 5. For different delays even large delays, the control law can stabilize in the end. It can be seen that under the same parameters and conditions, a larger delay upper bound leads to slower convergence speed, but has little effect on the steady-state errors. The proposed control law is robust to communication delays. At a steady state, the control law can reduce the errors to a low level. However, as the communication delay increases, the information received by each robot becomes more delayed, which can affect the dynamic performance.
Figure 6: A comparison of the convergence speed with other methods.
Figure 7: A comparison of the robustness with other methods.
Due to the difficulty in fully replicating existing approaches, the basic leader-follower control strategies similar to [20, 35] were conducted. The setup is consistent with Fig. 4. It is assumed that the time delay is zero. The first comparison is the convergence speed. The parameters in the control law are set as the same and the result (40s to 45s) is shown in Fig. 6. When the proposed control law is applied at 40s, it takes only 0.84s to reach stability. Applying other control strategies takes 2.52s to reach stability. The convergence speed improves by 66.67%. The second comparison is robustness. A large velocity disturbance is applied to one robot to test the impact on the control laws. Let the end-effector velocity of Robot 1 become [$0.1+\Omega(t)$, $0.1+\Omega(t)$]${}^{\text{T}}$m/s. The result (40s to 50s) is shown in Fig. 7. The average error of the proposed control law from 45s to 50s is 1.2724N, while the average error of other control strategies is 2.9547N. When the disturbance is large, the error is reduced by 56.94%. The comparison shows that the proposed control law has significant advantages in both convergence speed and robustness.
4 Experiments and Discussions
4.1 Experimental Setup
The proposed control law is verified in physical experiments. Each robot is composed of a HANGFA omnidirectional wheeled platform and a JAKA Zu7 manipulator. Each robot is equipped with a Robotiq FT 300-S Force Torque Sensor and a Robotiq 2F-85 Gripper as the end-effector, and an onboard Next Unit of Computing (NUC) computer with Intel Core i7-1165G7 CPU at 2.80 GHz (4 cores) and 16GB RAM. Each onboard computer independently calculates the motion of its respective robot according to the control law. The control frequency of the robots is 25 Hz. Besides, all robots are in the same local network for communication. The communication framework is implemented and achieved through ROS-melodic. Generally, the communication delays are less than 0.01s. The physical robot is shown in Fig. 8.
Figure 8: The physical model of a mobile manipulator robot used in experiments.
@@ 4.1 Experimental Setup
Figure 10: Comparison between applying and not applying the control law. The interaction wrenches in the world frame without (with) applying the control law is represented by the red (blue) curve.
Figure 11: Two mobile manipulator robots cooperatively manipulate an object to perform a reciprocating rotational motion.
Figure 12: Comparison between applying and not applying the control law. The interaction wrenches in the world frame without (with) applying the control law is represented by the red (blue) curve.
Figure 13: Two mobile manipulator robots cooperatively transport an object.
Figure 14: Comparison between applying and not applying the control law. The force-torque sensor data without (with) applying the control law is represented by the red (blue) curve.
4.2 Experimental Results
@@
The process of the experiment is illustrated in Fig. 9, and the motion duration is 83.2s. The experiment is repeated twice, once without the control law and once with. The results of interaction wrenches are shown in Fig. 10. Without the control law, the interaction wrenches could not be maintained at the desired values. Especially in the $x$-direction, the forces are very large due to the pushing and pulling between the robots when cooperatively manipulating the object. Additionally, in the $z$-direction, the robots could not evenly share the load. In contrast, with the control law applied, the interaction wrenches are maintained at the desired values in all six directions. Compared to the scenario without the control law, implementing the control law reduces the maximum errors by [93.37, 66.62, 92.38, 68.75, 58.82, 64.24]and the average errors by [98.68, 93.49, 96.58, 55.36, 61.96, 84.89]for robot1 in the corresponding directions, and the maximum errors by [94.38, 72.55, 93.30, 75.74, 70.18, 26.05]and the average errors by [98.56, 90.15, 97.30, 74.94, 73.36, 37.10]for robot2 in the corresponding directions. The tracking errors of the end-effector $||\bm{e}^{\text{d}}-\bm{e}||_{2}$ are shown in Fig. 10(c). Applying the control law reduces the average tracking errors for robot 1 and robot 2 by 79.62% and 82.93%, respectively. As previously mentioned, due to the tight contact between the robot and the object, even small errors can lead to significant interaction wrenches.
Further, we conduct experiments with a more rigid object, as shown in Fig. 11. The object is replaced with a wooden board weighing 80N. The corresponding results are shown in Fig. 12. It can be seen that the proposed control law is still effective in reducing interaction wrenches and maintaining them at the desired values. Applying the control law reduces the average tracking errors for robot 1 and robot 2 by 92.52% and 91.82%, respectively.
@@
The process of the experiment is illustrated in Fig. 13, and the motion duration is 130.4s. From 17s to 57s, the platforms move forward simultaneously to avoid exceeding the reachable space of the manipulators. From 70s to 100s, the platforms move laterally simultaneously to transport the object. As shown in Fig. 14, it can be seen that without the control law, the interaction wrenches vary greatly, especially in the $f_{x}$ directions. Compared to the scenario without the control law, implementing the control law reduces the maximum errors by [45.97, 54.18, 53.56, 81.85, 77.39, 21.52]and the average errors by [85.92, 62.72, 60.49, 93.13, 88.80, 67.99]for robot1 in the corresponding directions, and the maximum errors by [45.47, 34.97, 52.53, 73.38, 74.41, 46.05]and the average errors by [86.38, 58.44, 69.38, 85.51, 83.32, 72.44]for robot2 in the corresponding directions. The tracking errors of the end-effector are shown in Fig. 10(c). Applying the control law reduces the average tracking errors for robot 1 and robot 2 by 91.39% and 92.15%, respectively.
@@
4.3 Discussion and Limitation
The stability of the proposed control law has been proven, and subsequent simulations and experiments have demonstrated that the proposed control law effectively maintains the desired interaction wrenches. Due to various factors in physical robots, such as kinematic parameter uncertainties and external disturbances, deviations may occur between the actual and planned results. This can lead to excessive interaction wrenches in cooperative object manipulation. The proposed control law can be used to address these situations. Compared to other control methods, the proposed control law only utilizes local sensor information, and does not require accurate dynamic parameters or torque control. These simple conditions make the control law applicable in most scenarios and more suitable for real-world systems. Additionally, the communication delay between robots is considered, and it has been proven that the control law is robust against delays.
However, the proposed control strategy also has some limitations. First, the stiffness of the object is required. This parameter often requires empirical values or preliminary experimental determination. Improper setting of control coefficients may lead to instability or divergence of the control law. Second, since only the joint velocity is considered, the control law may fail in scenarios where the trajectory is demanding, such as when the velocity of the object changes frequently and largely.
5 Conclusion and Future Work
This paper aims to reduce the interaction wrenches during the cooperative object manipulation of multiple mobile manipulator robots. A distributed motion control law based on local information is proposed. The stability of the control law has been rigorously proven. The effects of graph connectivity and communication delays on the control law have been studied in simulations. The effectiveness of the control law has been validated in physical experiments. In conclusion, the proposed control law is easily implemented in physical systems and general to different robots.
Regarding future work, on the one hand, we will further refine the control strategy. The dynamics model of the robot will be considered to reduce impact forces when the robot’s velocity changes. On the other hand, we will study the rigorous proof of the control law’s robustness to disturbances and parameter uncertainties.
[…]